Celsius temperatures readings can be converted to farenheit readings using the fourmula f= 9/5 C + 32. What is fareheit temperature that corresponds to each of the following celsius temperatures: -10, 0, 15, 100?
Answer by doukungfoo(195)

What you need to do is use the formula given:
f= 9/5 C + 32
For each celsius temperature given just plug in that value for C in the equation.
Example:
What is the fareheit temperature for the celsius temperature -10 degrees?
f= 9/5 C + 32
f= 9/5 (-10) + 32
f= -90/5 + 32
f= -18 + 32
f= 14
So -10 degrees celsius corresponds to 14 degrees fareheit
Lets try another one. How about 0 degrees celsius?
f= 9/5 C + 32
f= 9/5 (0) + 32
f= 0 + 32
f= 32
So 0 degrees celsius corresponds to 32 degrees fareheit
Now you should be able to find the last two by following these examples.